OpenStudy (helpblahblahblah):

Help please How many solutions does the equation 6a − 3a − 6 = −2 + 3 have? A. Two B. None C. One D. Infinitely many

OpenStudy (helpblahblahblah):

I think B

OpenStudy (xapproachesinfinity):

why?

OpenStudy (helpblahblahblah):

Because I couldn't find a correct solution

OpenStudy (xapproachesinfinity):

how do you solve equation involving one variable?

OpenStudy (xapproachesinfinity):

well since you said let me ask how did you decide that you couldn't find any correct solution what did you try and made think there is not solution

OpenStudy (xapproachesinfinity):

you should have by now some ideas?

OpenStudy (xapproachesinfinity):

sorry of the english lol missing lot of words, but you got me i hope

OpenStudy (helpblahblahblah):

I'm kinda lost tbh lol...

OpenStudy (xapproachesinfinity):

i need to see you work something whether be wrong or not but tell me whatever you can do here to solve such equation

OpenStudy (helpblahblahblah):

6a-3a-6=-2+3 Combined liked terms: 3a-6=1 Added 6 on each side left me with 3a=7

OpenStudy (xapproachesinfinity):

alright great! actually you know how to do this i'm asking myself why are you saying you lost?

OpenStudy (xapproachesinfinity):

now you got 3a=7 what else can you do?

OpenStudy (helpblahblahblah):

Divide 3 on each side

OpenStudy (xapproachesinfinity):

good so a=...

OpenStudy (helpblahblahblah):

2.3?

OpenStudy (xapproachesinfinity):

yes but let's just leave 7/3 no need for decimal in this case unless asked don't do

OpenStudy (xapproachesinfinity):

so how many solution is that?

OpenStudy (helpblahblahblah):

One?

OpenStudy (xapproachesinfinity):

yes! see this is nothing really, just some simple algebra

OpenStudy (helpblahblahblah):

Thanks :)

OpenStudy (xapproachesinfinity):

apparently you are good at this, you just lack the motivation

OpenStudy (xapproachesinfinity):

and no problem :)

OpenStudy (helpblahblahblah):

:)

OpenStudy (anonymous):

:(

OpenStudy (helpblahblahblah):

Why that face? @jacobrocket173

OpenStudy (anonymous):

fan me

OpenStudy (helpblahblahblah):

I did like just now

OpenStudy (acxbox22):

fans are earned bro

OpenStudy (helpblahblahblah):

Oh no well I know him I just hadn't fanned him again @acxbox22